Copy equipment was acquired at the beginning of the year at a cost of $56,000 that...

  1. Copy equipment was acquired at the beginning of the year at a cost of $56,000 that has an estimated residual value of $8,000 and an estimated useful life of 5 years.   It is estimated that the machine has an estimated 1,000,000 copies.  This year 240,000 copies were made.  Determine the (a) depreciable cost, (b) depreciation rate, and (c) the units-of-production depreciation for the year.
Answer #1

a) depriciation cost = 56000 - 8000 =48000


b) depriciation rate = 48000/1000000 =0.48


c)units of production depriciation = 0.48 * 240000 =115200